40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 126 of 346  Not logged in ELOG logo
ID Date Author Type Category Subject
  11097   Wed Mar 4 03:42:14 2015 JenneUpdateLSCTried a few CARM / ALS fool things, no success

Much of tonight was spent fighting with ETMX.  This time, ASC was definitely off, there was nothing coming out of the ASC filter banks except the static output of the ASS.  I tried turning off the 1000 count POS offset, but I think that made it a little worse. I ended up putting the offset back.

It's a little confusing, since it sometimes moves when there is no LSC actuation.  However, it definitely moves when there is some LSC actuation.  I did a test where every time I enabled the IR arm locking and caught lock, I saw a step in the SUSPIT and SUSYAW error signals.  Once lock was aquired, it would settle and stay somewhere. If I unlocked the cavity, there was no "undo" step - it just stayed where it was.  I wasn't letting it sit long enough to see if it spontaneously moved during this test.

Here's a plot of this test.  The only button I'm touching is the LSC enable button.  ASC is off, ASS is frozen (DC values exist, but no dither, no feedback).  This was done when the 1000 count POS offset was off. The steps are less bad when the offset is on.

Inline image 1


In between fighting with the ETM, I was able to do several trials with the PRFPMI. 

I was playing with CARM and ALS fool.

First, I used REFL55 normalized by the sum of the transmissions as the error signal for the MC filter bank and saw that REFL11 (as an out of loop signal) got much more smooth, and centered around zero.  However, I wasn't able to get the same thing with REFL11.  No matter the sign I used for the MC filter bank, the IFO would squeak (some high freq gain peaking I think), and then I'd lose lock.  This was true whether I used REFL11 through the common mode board or just directly into the ADC.


Just now, I did one trial of switching DARM over to AS55Q, just to grab a spectra of the MICH noise that Q and I saw yesterday.

PRFPMI_DARMonAS55Q_3Mar2015.pdf


I'm a little confused by some delay that seems to exist between the "A" and "B" error signals (right after the LSC input matrix) and the _IN1 point of the servo filters.  I didn't save the measurement (bad Jenne), but there's a ~40 degree difference between DARM_A_ERR/DARM_IN2 and DARM_IN1/DARM_IN2.  I don't think there should be anything there.  Anyhow, it makes the DARM loop measurements look funny.  If you just look at, say, DARM_B_ERR/DARM_IN2, you'll think that there's no way that the loop will be stable.  However, it will actually be fine. 

For tomorrow, we should take the DARM loop measurement with much less actuation.  As with last night, I blew the lock by trying to measure the DARM loop.

Attachment 1: PRFPMI_DARMonAS55Q_3Mar2015.pdf
PRFPMI_DARMonAS55Q_3Mar2015.pdf
  11096   Wed Mar 4 00:50:36 2015 ranaBureaucracyTreasureTransitioned DARM to AS55Q, some other work

Just in case there was some confusion, the champagne on my desk is not to be opened before I get back, no matter how many signals are transitioned to RF.

  11095   Tue Mar 3 19:19:54 2015 JenneUpdateLSCTransitioned DARM to AS55Q, some other work

[Jenne, EricQ]

A slightly more coherent elog for last night's work.


All night, we've been using REFL165 to hold the PRMI.  It's working very nicely.  To help it catch lock, I've set the gain in the PRCL filter bank high, and then the *0.6 filter triggers on.  The carm_cm_up script now will lock the PRMI on REFL165. 

We had to reset the REFL165 phase after we acquired lock - it was -91, but now is -48.  I'm not sure why it changed so significantly from the PRMI-only config to the PRFPMI config.


We measured the ALS fool cancellation with the arms held off resonance, at arm powers of a few.  Although, they were moving around a lot, but the measurement stayed nice and smooth.  Anyhow, we get almost as good of cancellation as we saw with the single arm (after we made sure that both phase trackers had the same UGF):

ALSfool_PRFPMI_2Mar2015.pdf

We were able to partly engage it one time, but we lost lock at some point.  Since the frame builder / daqd decided that that would be just the *perfect* time to crash and restart, we don't have any frame data for this time.  We can see up to a few seconds before the lockloss, while we were ramping up the RF PD loop gain though, and MICH was hitting the rails.  I'm not sure if that's what caused the lockloss, but it probably didn't help.

The ALS fool gain was 22, and we were using FMs 4, 6 (the pendulum and Rana's "comp1"), the same filters that were used for the single arm case.  The LSC-MC filter bank gain lost lock when we got to about 5.6 (we were taking +3dB steps).

We were using REFL55I/(TRX + TRY) as our CARM RF error signal.  We were using REFL55 rather than REFL11 because we were worried that REFL11 didn't look good - maybe it was saturating or something.  To be looked into.

Here's the striptool that was running at the time, since we don't have frame data:


At this point, since we weren't sure what the final gain should be for the RF CARM signal, and we could sit at nice high arm powers (arm powers of 30ish correspond to CARM offsets of about 50pm), we decided to try just a straight jump over to the RF signals. 

The first time around, we jumped CARM to (-0.2)*REFL55/(TRX+TRY), but we only stayed lock for 1 or 2 seconds.  That was around 1:55am.

We decided that perhaps it would be good to get DARM moved over first, since it has a much wider linewidth, so the rest of the trials for the night were transitioning DARM over to (0.0006)*AS55Q/(TRX + TRY).  AS55 was saturating, so we reduced its analog gain from 18dB to 9dB and re-ran the LSC offsets.  The MICH noise was pretty high when we were at low CARM offset, although we noticed it more when DARM was on AS55.  In particular, there is some peak just below 40Hz that is causing a whole comb of harmonics, and dominating the MICH, PRC and DARM spectra.  I will try to get a snapshot of that tonight - I don't think we saved any spectra from last night.  Turning off DARM's FM3 boost helped lower the MICH noise, so we think that the problem is significant coupling between the two degrees of freedom. 

After the first one or two tries of getting DARM to AS55, we started engaging the arm ASC loops - they helped the dark port contrast considerably.  The POP spot still moves around, but the dark port gets much darker, and is more symmetric with the ASC on.

Attachment 1: ALSfool_PRFPMI_2Mar2015.pdf
ALSfool_PRFPMI_2Mar2015.pdf
Attachment 2: ALSfool_kindaEngaged_2March2015_noFrames.png
ALSfool_kindaEngaged_2March2015_noFrames.png
  11094   Tue Mar 3 19:19:15 2015 ericqUpdateIOOPC Drive / FSS Slow correlation

Jenne and I were musing the other night that the PC drive RMS may have a "favorite" laser temperature, as controlled by the FSS Slow servo; maybe around 0.2.

I downloaded the past 30 days of mean minute trend data for MC Trans, FSS Slow and PC Drive, and took the subset of data points where transmission was more than 15k, and the FSS slow output was within 1 count of zero. (This was to exclude some outliers when it ran away to 3 for some days). This was about 76% of the data. I then made some 2D histograms, to try and suss out any correlations. 

Indeed, the FSS slow servo does like to hang out around 0.2, but this does not seem to correlate with better MC transmission nor lower PC drive.

In the following grid of plots, the diagonal plots are the 1D histograms of each variable in the selected time period. The off diagnoal elements are the 2D histograms. They're all pretty blob-y, with no clear correlation. 

Attachment 1: jointplot.png
jointplot.png
  11093   Tue Mar 3 11:38:11 2015 SteveUpdatesafetyprofessional crane inspection

 

Quote:

Safety glasses were measured and they are all good. I'd like to measure your personal glass if it is not on this picture.

Quote:

Safety audit went soothly. We thank all participients.

Correction list:

1, Bathroom water heater cable to be stress releived and connector replaced by twister lock type.

2, Floor cable bridge at the vacuum rack to be replaced. It is cracked.

3, Sprinkler head to be moved eastward 2 ft in room 101

4, Annual crane inspection is scheduled for 8am Marc 3, 2015

5, Annual safety glasses cleaning and transmission measurement will get done tomorrow morning.

 

Konecranes' Fred inspected and load tested all tree cranes at with 450 lbs

Attachment 1: 450lbs2015.jpg
450lbs2015.jpg
  11092   Tue Mar 3 09:42:40 2015 SteveUpdateSUSETMX is drifting

The temperature of the east and south ends  are normal, they are about the same.

 

Attachment 1: ETMXdrifting.png
ETMXdrifting.png
Attachment 2: ETMXvsTemp.png
ETMXvsTemp.png
  11091   Tue Mar 3 04:43:12 2015 JenneOmnistructurePEMOttavia is dead?

After some searching, including help from 4 security guys (I think they don't have a lot to do at 4:30am :), we found that Ottavia is super warm, and smelled burn-y.  She has been powered down and unplugged.  Security guys may call Steve's desk to follow up later today.sad

  11090   Tue Mar 3 04:41:45 2015 JenneUpdateLSCTransitioned DARM to AS55Q, some other work

Better elog tomorrow - notes for now:

REFL165 for PRMI has been "a champ" (quote from Q).  We're able to sit on ALS at average arm powers of 30ish.  Nice. 

Some ALSfool work - measured cancellation almost as good as single arm. 

One time transitioned CARM -> normalized REFL55I

Many times did DARM -> normalized AS55Q, see lots of noise at 39ish Hz - may be coupling from MICH??

Arm ASC loops helped improve dark port contrast. 

Note to selfs:  Need to make sure DTT templates have correct freq ordering - must be small freqs to large freqs.

  11089   Tue Mar 3 02:43:06 2015 JenneOmnistructurePEMNo heat??

It's super cold in the control room and EE bench area tonight.  I'm wondering if, similar to what happened on Dec 29th (http://nodus.ligo.caltech.edu:8080/40m/10846) the campus steam is off?  Or just our heater is broken?  The thermostat is cranked up to 80 over by the bathrooms (this is usually ~74F), but we're still cold.

It's 69F in the control room right now (usually mid-high 70s).

EDIT, JCD, 4am: It's 64.3F in the desk area, 67.8F in the control room.  It also smells in the control room like some heater has been off for a while, and is turning back on - that burned dust smell that happens after you haven't turned on the heater all summer. 

EDIT again: The burn-y smell is getting stronger I think.  Security is sending someone over to come check it out.

  11088   Mon Mar 2 18:34:12 2015 JenneUpdateSUSETMX moving during single arm lock

This has been edited several times over the last several hours, as I try to change different parameters, to see if they affect the movement of ETMX.  So far, I don't know what is causing the motion.  If it is there, it is only present when the LSC is engaged, so I don't think it's wobbling constantly on a twisted wire.

FINAL EDIT, 9:10pm:  The arm ASC was turning itself on when the arms were locked.  Whelp, that was only 3 hours of confusion.  Blargh.

For his penance for leaving the arm ASC engaged, Q has made a set of warning lights on the LSC screen, right next to the ASS warning lights.


ETMX might be having one of those days today, which is lame. 

So far tonight, I have run the LSC offset script, set the FSS slow value to +0.2, and run the arm ASS scripts.  Nothing too crazy I think.

Sometimes when I lock the single arms, the ETMs move around like crazy.  Other times, not.  What is going on here???  The ETMs don't move at all when they are not being actuated on with the LSC.

In this screenshot you can see the end of a POX/POY lock stretch where everything was nice and good.  Then, the arms were unlocked, and they have a bit of a DC offset.  After settling from that step, they continue sitting nice and still.  Then, I relock the cavities on POX and POY a little before -4 minutes.  ETMY takes a moment to pull itself together, but then it's steady.  ETMX just wobbles around for several minutes, until I turn off the LSC enable switch (happened after the end of this plot). 

I'm not going to be able to lock like this.  Eeek! 


This is somehow related to light being in the Xarm.  This next plot was taken while the arms were held with ALS in CARM/DARM mode. 

  • Between -4 and -3, and again between -2 and -1.5, the arms were held with ALS "on resoanance".  Yarm power (green) is fine, Xarm power (purple) is not fine.  The ETMX OSEM pitch and yaw signals are bumpy at the same times.
  • Between -3 and -2 minutes, the arms were held with ALS off resonance.  Obviously you can't see anything about IR transmission, but you can see that the SUSPIT and SUSYAW signals quiet down. 
  • Between -1 and 0, I moved the Yarm ALS offset such that only it was resonating (the jags at the beginning of the resonance is me finding it, not to do with ETMY misbehaving). 
    • The ETMX SUS signals stay quiet. 
    • ETMY YAW has a DC offset when there is IR power in the arm.
  • After 0 minutes on this scale, I move the offsets such that only Xarm is "on resonance". 
    • The ETMX osem signals go bumpy again.


I closed and re-opened all 3 green shutters.  Now (at least the last 8 arm locks in the last 6 mintues) ETMX has never gone wobbly, except for a little bit right after acquisition, to deal with whatever the DC offset it. Why is this changing? 


The arms were fine for one long ~30 minute lock while I stepped out for dinner.  At some point after returned, the MC lost lock. When the arms came back, ETMX was being fussy again.  Then, it decided that it was done.

In this plot, at -1 minute I started the ASS.  Other than that, I did not touch any buttons at all, just observed.  I have no idea why at about -3 minutes the bad stuff seems to go away. 

I was curious if it had to do with the DC pointing of the optics, so I unlocked the arms, put ETMX about where it was during the long good lock stretch, then reaquired lock.  I had to undo a little of that so that it would lock on TEM00, but at the beginning of the lock stretch (starting at about -3) the pitch is about the same spot.  But, the oscillations persist.  This time it was clear that the oscillations were around 80 mHz, and they started getting bigger until they settled to an amplitude they seemed to like.


Seems pretty independent from FSS temp.  There are 3 lock stretches in the next plot (easier to see by looking at the Yarm transmission, green trace).  The first one, the FSS slow was at 0.35.  the middle one, it was around 0.05.  The last one, it was around -0.4.  Other than the different DC pointings (which I don't know if they are related), I don't see anything qualitatively different in the movement of ETMX.

Attachment 1: ETMX_moving.png
ETMX_moving.png
Attachment 2: ETMXmoving_ALSlock.png
ETMXmoving_ALSlock.png
Attachment 3: StartsMoving_stopsItself_ranASS.png
StartsMoving_stopsItself_ranASS.png
Attachment 4: 80mHz.png
80mHz.png
Attachment 5: IndependFromFSStemp.png
IndependFromFSStemp.png
  11087   Mon Mar 2 17:02:01 2015 ericqUpdateLSCBS - PRM decoupling

Using PRX, I remeasured the relative actuation strengths of the BS and PRM to see if the PRM correction coefficient we're using is good. 

My result is that we should be using MICH -> -0.2655 x PRM + 0.5*BS.

This is very close to our current value of -0.2625 x PRM, so I don't think it will really change anything.


Measurement details:

The reason that the BS needs to be compensated is that it really just changes the PRM->ITMX distance, lx, while leaving the PRM-ITMY distance, ly, alone. I confirmed this by locking PRY and seeing no effect on the error signal, no matter how hard I drove the BS. 

I then locked PRX, and drove an 804Hz oscillation on the BS and PRM in turn, and averaged the resultant peak heights. I then tried to cancel the signal by sending the excitation with opposite signs to each mirror, according to their relative meaured strength.

In this way, I was able to get 23dB of cancellation by driving 1.0 x PRM - 0.9416 * BS. 

Now, in the PRMI case, we don't want to fully decouple like this, because this kind of cancellation just leaves lx invarient, when really, we want MICH to move (lx-ly) and PRCL to move (lx+ly). So, we use half of the PRM cancellation to cancel half of the lx motion, and introduce that half motion to ly, making a good MICH signal. Thus, the right ratio is 0.5*(1.0/0.9416) = 0.531. Then, since we use BS x 0.5, we divide by two again to get 0.2655. Et voila.

  11086   Fri Feb 27 21:16:10 2015 JenneUpdateLSCPRMI locked during CARM scan

Here are a bunch of PDFs of time series from last night's CARM sweeps.  The y-axes are all calibrated (except for the TRX/TRY, which are just normalized to single arm power, as usual) to real units - meters for the error signals, and Newtons for the control signals.  The y-axes for each plot are the same on all PDFs (ex, the control signal plot in the lower left has the same range for all cases) so that it is easy to compare directly. 

The most striking thing is that while the PRMI is held on REFL33, the MICH control signal saturates as we go through arm resonance.  If the PRMI is held on REFL165, there is no such problem.  I think we're going to have a lot more luck keeping the PRMI on REFL 165.


Plots while held on REFL 33:

REFL33lock_431.pdf

REFL33lock_516.pdf

REFL33lock_605.pdf

REFL33lock_648.pdf


Plots while held on REFL 165:

REFL165lock_835.pdf

REFL165lock_877.pdf

REFL165lock_912.pdf

REFL165lock_940.pdf

 

Attachment 1: REFL33lock_431.pdf
REFL33lock_431.pdf
Attachment 2: REFL33lock_516.pdf
REFL33lock_516.pdf
Attachment 3: REFL33lock_605.pdf
REFL33lock_605.pdf
Attachment 4: REFL33lock_648.pdf
REFL33lock_648.pdf
Attachment 5: REFL165lock_835.pdf
REFL165lock_835.pdf
Attachment 6: REFL165lock_877.pdf
REFL165lock_877.pdf
Attachment 7: REFL165lock_912.pdf
REFL165lock_912.pdf
Attachment 8: REFL165lock_940.pdf
REFL165lock_940.pdf
  11085   Fri Feb 27 14:17:51 2015 SteveUpdatesafetysafety audit 2015

Safety glasses were measured and they are all good. I'd like to measure your personal glass if it is not on this picture.

Quote:

Safety audit went soothly. We thank all participients.

Correction list:

1, Bathroom water heater cable to be stress releived and connector replaced by twister lock type.

2, Floor cable bridge at the vacuum rack to be replaced. It is cracked.

3, Sprinkler head to be moved eastward 2 ft in room 101

4, Annual crane inspection is scheduled for 8am Marc 3, 2015

5, Annual safety glasses cleaning and transmission measurement will get done tomorrow morning.

 

Attachment 1: 1064Tchecked.jpg
1064Tchecked.jpg
  11084   Fri Feb 27 11:20:49 2015 ericqUpdateComputer Scripts / ProgramsiPython Notebook for LSC Sensing Matrix
Quote:

** along the way, I noticed that the reason this notebook hasn't been working since last night is that someone sadly installed a new anaconda python distro today  without telling anyone by ELOG. This new distro didn't have all the packages of the previous one.no I've updated it with astropy and uncertainties packages.

My bad, sorry! 

Yesterday, I was trying to install a package with anaconda's package manager, conda, but it was crashing in some weird way. I wasn't able to fix it, which led me to create a fresh installation. 

  11083   Fri Feb 27 10:25:24 2015 steveUpdateCamerasquad picture problem

MUX input 7 to ITMXF camera cable was replaced by temporary cable labeled as 888

The problem remains to be the same black stripe at the bottom of the image The single picture is OK.

Attachment 1: quadMUX.jpg
quadMUX.jpg
  11082   Fri Feb 27 03:48:29 2015 JenneUpdateLSCPRMI locked during CARM scan

One of the things that I looked at tonight was whether or not I could hold the PRMI on REFL165 at CARM offset of 0, and it turns out that I can.  Hooray.  The next step was having a look to see if it is actually less noisy than the REFL33 lock.

I calibrated REFL33 and REFL165 to meters (I have the data to do the same for 11 and 55, but haven't done so yet).  This way, we can directly compare the signals from each PD. 

PD calibrations to meters [m/ct]
REFL 33 I 3.2e-11
REFL 33 Q 7.3e-12
REFL 165 I 9.5e-12
REFL 165 Q 1.2e-11

I scanned between +3 and -3 CARM digital offset (which we think is about 1nm/count while held on ALS), with a ramp time of 10 seconds.  I did this several times while the PRMI was locked on both REFL33 and REFL165. Here are the gps times for 8 examples where the PRMI did not lose lock during the sweep:

PD for lock CARM sweep GPS start time
REFL33 +3 -> -3 1109064431
REFL33 +3 -> -3 1109064516
REFL33 -3 -> +3 1109064614 (ended)
REFL33 -3 -> +3 1109064648
REFL165 +3 -> -3 1109064835
REFL165 -3 -> +3 1109064877
REFL165 +3 -> -3 1109064912
REFL165 -3 -> +3 1109064940

Here are screen shots from the first REFL33 sweep, and the first REFL165 sweep. DTT can't print 3 plots together, so I'll have to make this nicer later.  The top plot is the error signals, calibrated to meters.  The middle plot is the control signals, that need to be calibrated to Newtons.  The bottom plot is the arm powers, so you can see roughly where we were in the sweep.

We'd like to see a MIST simulation, or perhaps e2e, to see what the predicted disturbance is for each of the error signals during the CARM resonance.  We want to make sure that the loops are engaged for all of the degrees of freedom for the simulation.


Recipes for tonight:

REFL165 sometimes has a tough time catching lock by itself, but if you add either REFL33 or REFL55 error signals to the REFL165 signals, it'll catch, and then you can just remove the extra error signals.  Also, it doesn't stay locked very robustly unless you include the PRCL FM1 boost.

PRCL in PRFPMI, REFL55 I&Q MICH PRCL
Input Matrix 60* REFL55Q 6* REFL55I
Gain 3 -0.02
DoF trigger POP22I + 0.1*POPDC, 50 up 0.5 down POP22I + 0.1*POPDC, 50 up 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 2; 50 up, 0.5 down, 5 sec wait FM 1, 2, 9; 50 up, 0.5 down, 1 sec wait
Normalization none none
Output matrix 0.5*BS - 0.2625*PRM 1*PRM

 

PRCL in PRFPMI, REFL165 I&Q MICH PRCL
Input Matrix -0.6* REFL165Q 2.2* REFL165I
Gain 3 -0.02
DoF trigger POP22I + 0.1*POPDC, 50 up 0.5 down POP22I + 0.1*POPDC, 50 up 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 2; 50 up, 0.5 down, 5 sec wait FM 1, 2, 9; 50 up, 0.5 down, 1 sec wait
Normalization none none
Output matrix 0.5*BS - 0.2625*PRM 1*PRM
Attachment 1: REFL33Lock.png
REFL33Lock.png
Attachment 2: REFL165Lock.png
REFL165Lock.png
  11081   Fri Feb 27 01:59:57 2015 ranaHowToLSCiPython Notebook for LSC Sensing Matrix

I have adapted one of Evan's python scripts into an ipython notebook for calculating our PRMI sensing matrix - the work is ~half done.

The script gets the data from the various PD channels (like REFL33_I) and demdoulates it at the modulation frequencies. At the moment its using just the sensing channels, but with the recent addition of the SUS-LSC_OUT_DQ channels, we can demod the actuation channels as well and not have to hand code the exc amplitudes and the basolute phase. Please ignore the phase for the moment.

The attached PDF shows the demod (including lowpass) outputs for a 2 minute stretch of PRMI locked on f2. Next step is to average these numbers and make the radar plots with the error bars. The script is scripts/LSC/SensingMatrix/PRMIsensMat.ipynb and is in the SVN now.

** along the way, I noticed that the reason this notebook hasn't been working since last night is that someone sadly installed a new anaconda python distro today  without telling anyone by ELOG. This new distro didn't have all the packages of the previous one.no I've updated it with astropy and uncertainties packages.

I've fixed the Radar plot making part, so that's now included too. The radial direction is linear, so you can see from the smearing of the blobs that the uncertainty is represented in the graphics due to each measurement being a small semi-transparent dot. Next, we'll put the output of the statistics on the plot: mean, std, and kurtosis.

Attachment 1: Plots_1109056456.pdf
Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf Plots_1109056456.pdf
Attachment 2: Radar_1109056456.pdf
Radar_1109056456.pdf
  11080   Thu Feb 26 20:02:46 2015 JenneUpdateLSCre: some thoughts

I have clarified my elog from last night to indicate that the sensing matrix in the "33MHz cancellation" configuration was measured with the PRMI held on REFL55 I&Q.

Also, I just re-read my control room notes from yesterday, and I typed the wrong demod phases into the table last night.  The elog has been edited.  Most significantly, the REFL33 demod phase did not change by 70 degrees.  It did change by 10 degrees, but that is likely from the fact that it was formerly tuned for PRCL in PRFPMI, and last night I tuned it for PRCL in PRMI-only.

  11079   Thu Feb 26 16:46:37 2015 manasaUpdateGeneralIOO rack power supply

I shutdown the +/-15V and the +/-24V sorensons on the IOO rack to connect the FOL RF PDs to the rack power supply.

They were turned back ON after the work. PMC and MC were relocked.

Quote:

[Steve, Manasa]

We ran power cables and sma cables for the FOL fiber module from the PSL table to the IOO rack.

 

  11078   Thu Feb 26 16:37:21 2015 manasaFrogsLSCRIP illegal power supply

No more illegal power supply at the LSC rack yes

The amplifiers are now being powered by the rack power supply through fuse blocks.

To make new connections, I shutdown the +/-15 V low noise power supplies. They were turned back ON after the work.

Quote:

 Is this your illegally installed HP bench power supply?

20131023_222351.jpg

If so, or if not but you care about the signal that passes through these amplifiers, I suggest you remove this temporary power supply and wire the power from the rack power supplies through the fuse blocks and possibly use a voltage regulator.

In 24 hours, that power supply will be disconnected and the wires snipped if they are still there.

 

  11077   Thu Feb 26 13:55:59 2015 jamieUpdateComputer Scripts / ProgramsFB IO load
We should use "ionice" to throttle the rsync. Use something like "ionice -c 3 rsync ..." to set the priority such that the rsync process will only work when there is no other IO contention. See "man ionice" for other options.
  11076   Thu Feb 26 13:17:31 2015 ericqUpdateComputer Scripts / ProgramsFB IO load

Over the past few days, I've occasionally been peeking at the framebuilder IO load to see If I could correlate anything with it, but it's usually been low when I looked. I.e. with daqd and all models running, the %wa time was in the few percents at most.

Just now, I was seeing some EPICS sluggishness, and sure enough, the %wa was in the 50-60 range. I used iostat -xmh 5 on the framebuilder to see that /dev/sda, the /frames drive, was at 100% utilization, which means it was reading and writing as fast as it possibliy could. 

I ssh'd over to nodus, and with iotop found that an rsync job was running (rsync -am --exclude .*.gwf full 131.215.114.19::40m/full), and its IO rates corresponded very closely to the data read rates on the framebuilder from /frames. 

I killed the rsync process on nodus, and the %wa time on the framebuilder dropped to near zero. The ASS striptools, where I had noticed the sluggishness, immediately started updating faster.

While rsync is supposed to play nice with a system's IO demands, maybe it only knows about nodus's IO usage, not fb which is the underlying NFS server where the frames live. I think it would be good to throttle the bandwidth of these jobs to a specific bandwidth. 50MB/s seemed like too much, so maybe 10MB/s is ok?

  11075   Thu Feb 26 10:56:34 2015 manasaUpdateGeneralRunning cables

[Steve, Manasa]

We ran power cables and sma cables for the FOL fiber module from the PSL table to the IOO rack.

  11074   Thu Feb 26 01:53:35 2015 KojiUpdateLSCModelled effect of relative modulation phase

Ok... This is what I was afraid of, and it seems true.
i.e. the relation ship of the modulations for the 3f cancellation is making the PRCL signals cancel each other.

It agrees with Anamaria's analysis that 11x44 is the strongest component in aLIGO 27MHz signal.
In fact,

00x33 has the order of \frac{m_1^2 m_2}{16}+\frac{m_1^3}{48}

11x22 has the order of \frac{m_1^3}{16}

11x44 has the order of \frac{m_1^2 m_2}{8}

22x55 has the order of \frac{m_1^2 m_2}{16}

Therefore 11x44 is inherently the strongest contribution at 33MHz.
(And then, of couse, the signal amplitudes have additional dependences on the reflectivity
and the gain of the IFO at each freq)

If we believe this result, it may be difficult to exploit the benefit of the signals under the 3f cancellation.
We probably have to go back to the original idea of cancelling the 3f modulation by adding 3f modulation.
(i.e. Produce 33MHz signal by freq tripling, add this signal to Kiwamu's box to elliminate 3f.)

  11073   Thu Feb 26 01:51:39 2015 ericqUpdateLSCSideband HOMs

So, my previous post suggested that 44*11 products might be the dominating signals in our "nominal" setup. I suppose that this could be not so bad, since it's not too unlike -11*22; the 11MHz field couples into the PRC and reflects with a rapidly changing phase around PRC resonance, and 44MHz is antiresonant, so it is a good local oscillator for REFL33. 

However, it then occured to me that my previous HOM analysis only looked at the 11MHz and 55MHz sidebands. 

When extending this to all of the sidebands within 55MHz, I discovered a troubling fact:

With the IFO parameters I have, the second spatial order +22MHz and fourth spatial order +44MHz fields almost exactly coresonate with the carrier in the PRFPMI! (DR, too)

If this is true, then any REFL33 signal seems to be in danger if we have an appreciable amount of these modes from, say, imperfect modematching.

On the other hand, we've been able to hold the PRMI with REFL33 when ALS is "on resonance," so I'm not sure what to think. (As a reminder, this analysis is done with analytic formulae for the complex reflectivities of the arm cavities and coupled recycling cavities as a function of CARM, spatial order and field frequency. Source is attached.)

It seems the Y arm geometry is to blame for this.

Maybe we should try to measure/confirm the Y arm g-factor...

Attachment 1: C1_HOMcurves_PR.png
C1_HOMcurves_PR.png
Attachment 2: C1_HOMcurves_Y.png
C1_HOMcurves_Y.png
Attachment 3: C1_HOMcurves_X.png
C1_HOMcurves_X.png
Attachment 4: C1_HOMlist.zip
  11072   Thu Feb 26 00:20:54 2015 ericqUpdateLSCModelled effect of relative modulation phase

I'm working on some more modelling investigations of this whole situation. The main thing I wanted to do was to look at the complex field amplitudes / IFO reflectivity to see how the PDH signal is affected by different field components. 

I still have plenty more to do, but I got a result which I though I should share. In addition to Jenne's simulation, I also see that between our "nominal" and "canceled" states as defined in Kojis ELOG 11036, there is a factor of ~20 difference in the PRCL signal in REFL33. 

The plots below are kind of like "PDH Signal Budgets" of the two states. 

Specifically, the reason our gain gets reduced is that, in the "canceled" state, the 44*11 and 55*22 products conspire to weaken the signal by having a slope opposite to the -11*22 type products. In contrast, in our "nominal" case, all of the products slope together. 

However, this also predicts that the nominal REFL33 is more sensitive to Carrier*33 than to the signal we desire, -11*22. The only reason it ever worked seems to be the biggest contriubutor, the unexpected 44*11! 

The "residual" trace is the difference of REFL33 and the sum of the field products shown, to justify that all relevant products had been included. 

The simulation that produced this was set up to create 4 orders of modulation at each EOM, with 3 orders of sidebands on sidebands. The demodulation phase was taken by lining up a PRM excitation entirely along I, as we would do on the actual instrument. 

MIST Simulation files attached!

Attachment 1: 33budget_canceled.png
33budget_canceled.png
Attachment 2: 33budget_nominal.png
33budget_nominal.png
Attachment 3: 2015-02-ModPhase.zip
  11071   Wed Feb 25 23:48:57 2015 ranaUpdateLSCsome thoughts
  • Comparing just the 2 cases with locking on 33, it seems that the 55 MHz gain has changed by 14 dB instead of the 10 dB that we expected. Is it that we need to measure the modulation drive change more carefully, or just that the PRMI was aligned differently?
  • The 165 signal changed by a factor of 60 (35 dB) which is more consistent with a ~12 dB change in Gamma2, so not so far off.
  • The fact that the whole sensing matrix increases in amplitude between 33 and 55 lock makes me think that either the alignment was very bad for the 33 lock, or that the 33 signals have a significant offset; if that's the case, then we should do as LLO and set the digital offsets in the 33 signals by locking first on 55.
  • How does the REFL33 demod phase change by 70 deg?
  11070   Wed Feb 25 20:00:39 2015 JenneUpdateLSCDelay line un-installed again - sensing matrix comparisons

I have measured the sensing matrix for the PRMI at the REFL photodiodes for both the nominal configuration and the 33MHz cancellation configuration.  The nominal configuration measurements do not compare well with those from November (http://nodus.ligo.caltech.edu:8080/40m/10701) which makes me unhappy no.  Both sets of nominal config reported below are from today, after tuning the demod phases and making sure the MICH and PRCL loops looked the same as yesterday (esp. overall gain).  The cancellation config data is from last night.

Note that the magnitude for each photodiode is referred to its own "PD counts".  Since the electronics are different for each PD, and that is not taken into account here, you cannot directly compare an element from one PD to an element from another PD.  What you can do (which is most of what we need right now) is compare all the different measurements for a single photodiode.

PRCL Sensing elements 33 MHz cancellation, REFL55 I&Q lock Nominal, REFL55I&Q lock Nominal, REFL33I&Q lock
  Mag [PD counts / m] Phase [deg] Mag [PD counts / m] Phase [deg] Mag[PD counts / m] Phase [deg]
REFL 11 3.0e13 0.05 4.2e13 0.1 2.9e13 0.1
REFL 33 1.1e11 3.4 1.8e12 0.25 1.2e12 0.3
REFL 55 1.8e10 2.1 1.4e11 4.8 9.3e10 4.1
REFL 165 7.0e9 15.5 6.3e11 0.5 4.2e11 0.9

 

MICH Sensing elements 33MHz cancellation, REFL55 I&Q lock Nominal, REFL55 I&Q lock Nominal, REFL33 I&Q lock
  Mag [PD counts / m] Phase [deg] Mag [PD counts / m] Phase [deg] Mag [PD counts / m] Phase [deg]
REFL 11 3.0e12 2.8 4.1e12 3.0 2.9e12 3.3
REFL 33 1.1e10 3.7 1.7e11 3.5 1.2e11 3.9
REFL 55 2.1e9 30.3 1.4e10 27.7 1.1e10 30.2
REFL165 6.4e8 24.4 6.3e10 21.0 4.5e10 23.0

So, what I'm apparently seeing is that the magnitudes of the sensing matrix signals that are made using 55MHz (i.e. everything but REFL11) change when we go into the cancellation configuration, but the phases of the sensing elements do not change significantly.  Also, I am apparently seeing that REFL11 and REFL33 only have about 3 degrees of separation between the MICH and PRCL signals no matter what configuration is used.  This doesn't make a lot of sense, since we know that we can lock robustly on REFL33I&Q (it's been sitting there happily as I write this elog), so it seems crazy that we could actually be so degenerate.  Also, at the bottom of the elog that I wrote in November 2014, I show a sensing matrix where both REFL11 and REFL33 have about 45 degrees of separation between the MICH and PRCL signals.

I don't think I'm doing anything too crazy here, particularly with the phase.  For a given PD and given DoF, I find the magnitude of the peaks of the I and Q signals, and just do atan2(Q-signal, I-signal)*180/pi, and those are the numbers that go in the phase columns above. 


Before taking my measurements, I tuned up the demod phases for the PRMI-only case.  I think REFL11 may have previously been tuned for CARM when the arms were held with ALS, but I don't really recall.  Anyhow, now all 4 REFL PDs are tuned for PRMI-only.

This was done while the PRMI was locked with REFL 55 I&Q. 

EDIT, 26Feb2015: Last night I mixed up the REFL11 and REFL33 new demod phases.  Bold are the corrected version.  Also, note that REFL33 was formerly tuned for PRCL in PRFPMI, which may be why it changed by ~10 degrees.

  Old demod phase [deg] New demod phase [deg]
REFL11 20 131.7 +/- 0.1   74.7 +/- 0.1
REFL33 142.2 74.7 +/- 0.1    131.7 +/- 0.1 
REFL55 19

15.3 +/- 0.3

REFL165 -172 -170.0 +/- 0.2

Here's the recipe for locking REFL 55 I&Q in the nominal modulation configuration.  It's the same as the REFL33 I&Q lock that I was using today, except that for the REFL33 version, the matrix elements are both unity.

PRMI REFL55 I&Q, nominal configuration MICH PRCL

Input Matrix

36*REFL55Q 12*REFL55I
Gain 5.0 -0.028
DoF trigger POP22I; 50 up, 0.5 down POP22I; 50 up, 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 2, 3, 6, 9; 50 up, 0.5 down, 5 second wait FM 1, 2, 6, 8, 9; 50 up, 0.5 down, 1 second wait
Normalization none none
Output matrix -1*ITMX, +1*ITMY 1*PRM

 

  11069   Wed Feb 25 14:51:13 2015 JenneUpdateLSCDelay line un-installed again

And now I've removed the delay line, and am in the process of reverting the demod phases, etc.

  11068   Wed Feb 25 14:35:32 2015 SteveUpdatesafetysafety audit 2015

Safety audit went soothly. We thank all participients.

Correction list:

1, Bathroom water heater cable to be stress releived and connector replaced by twister lock type.

2, Floor cable bridge at the vacuum rack to be replaced. It is cracked.

3, Sprinkler head to be moved eastward 2 ft in room 101

4, Annual crane inspection is scheduled for 8am Marc 3, 2015

5, Annual safety glasses cleaning and transmission measurement will get done tomorrow morning.

Attachment 1: safety2015audit.jpg
safety2015audit.jpg
  11067   Wed Feb 25 14:18:28 2015 ranaUpdateLSCDelay line re-installed, measurements round 2

The Schnupp asymmetry is definitely not an important parameter (no need for Koji to explode). It only serves to give us a bigger Q phase signal slope, but is not significant for the I phase signals.

The main anomaly is that the REFL33 optical gain (W/m) seems to have been reduced so much by the phase and amplitude adjustment of the 55 MHz modulation signal. One guess is that the true 3f signal is being made not by the (2*f1 - (-f1)) beat, but by some higher order beat. In addition to the usual RF fields produced by the 2 modulations, we must consider that the sidebands on sidebands produce intermodulation fields just after the EOM and so the fields with which we interrogate the PRMI are more complicated.

Jenne's Optickle calculation today should show us what the sensing matrix contribution is from each pair of fields, so that we can have a sensing matrix signal budget.

 

  11066   Wed Feb 25 12:16:27 2015 KojiUpdateLSCDelay line re-installed, measurements round 2

WHAT? WHAT? WHAT? It's obviously opposite.

If the reflectivity of the front mirror is fixed (=PRM reflectivity), the finesse increases when the reflectivity of the end
mirror (=Compond mirror reflectivity) increases. i.e. 11MHz has higher finesse, 55MHz has lower finesse.

{\cal F} = \frac{\pi \sqrt{r_{\rm PRM} r_{\rm COM}}}{1-r_{\rm PRM} r_{\rm COM}}

If the reflectivity of the front mirror is fixed, the amplitude gain of the cavity is higher when the reflectivity of the end mirror increases. i.e. 11MHz has higher gain, 55MHz has lower gain

g_{\rm PRM} = \frac{t_{\rm PRM}}{1-r_{\rm PRM} r_{\rm COM}}
 

Quote:

Our Schnupp asymmetry is small (3.9cm, IIRC), so the transmission of the 11MHz signal out the dark port is small.  This means that the finesse of the PRC for 11MHz isn't so huge.  On the other hand, since 55MHz is a higher frequency, the transmission out the dark port is larger and is a closer match to the PRM transmission, so the finesse of the PRC for 55MHz is higher. 

 

  11065   Wed Feb 25 11:01:05 2015 manasaUpdateComputer Scripts / ProgramsNew screen for FOL PID loop

Created a new medm screen C1ALS_FOL_PID.adl for FOL PID loop control in /medm/als/master/

This is not currently linked to the sitemap screen.

  11064   Wed Feb 25 04:59:47 2015 JenneUpdateLSCDelay line re-installed, measurements round 2

[Jenne, EricQ, Rana]

We spent this evening measuring and thinking about our 3f signals, and the effect of the modulation cancellation. 

I reinstalled the delay line box, and reduced the modulation depth of the 55MHz signal, so that we are in the state of modulation cancellation - there should be almost no power at 33MHz injected into the vacuum.  I carefully tuned the demod phase of the REFL 11, 33 and 55 MHz PDs, and locked the PRMI on REFL55 I&Q.  The signal in REFL 165 was very tiny, so as best as I could tell, the demod phase that Koji found last week was correct. 

Here is a little record of what the demodulation phases should be, for the "nominal" and "cancellation" configurations, so that we don't have to continually use the time machine.

  "Nominal" configuration 3f modulation cancellation
REFL 11 20 deg 76.0 deg (tuned to nearest 0.5 deg)
REFL 33 142.2 deg 120.3 deg (tuned to nearest 0.1 deg)
REFL 55 19 deg 173.0 deg (tuned to nearest 0.5 deg)
REFL 165 -172 deg 18 deg (same number as last week)
AS 55 -166.1 deg -111.1 deg (same number as last week)

Also, here is the locking recipe for REFL55 I&Q in the cancellation configuration:

PRMI, 3f cancellation, REFL55 I&Q MICH PRCL
Input Matrix acquire with -2*REFL55Q, then put matrix to -15*REFL55Q -12*REFL55I
Gain 3.0 -0.1
DoF trigger POP22I; 50 up, 0.5 down POP22I; 50 up, 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 1, 2, 3, 6, 9; 50 up, 0.5 down, 5 second wait FM 1, 2, 6, 9; 50 up, 0.5 down, 1 second wait
Normalization none none
Output matrix -1*ITMX, +1*ITMY 1*PRM

With this setup, we measured the sensing matrix.  MICH had an excitation at 370.123 Hz with 8,000 counts to -ITMX+ITMY (this is about 0.3nm for each ITM), and PRCL had an excitation at 404.123 Hz with 50 counts to PRM.  For tonight, here is a PDF of the peaks in DTT.  The data is saved in /users/Templates/LSC_error_spectra/SensMat_PRMI_24Feb2015.xml. 

SensMat_PRMI_24Feb2015.pdf


Rana has proposed to us an idea for why the REFL 33 signal should be dominated by the 55*22 contribution, rather than -11*22.  Eric is in the process of checking this out with a Mist model to see if it makes sense. Here's the gist:

Our Schnupp asymmetry is small (3.9cm, IIRC), so the transmission of the 11MHz signal out the dark port is small.  This means that the finesse of the PRC for 11MHz isn't so huge.  On the other hand, since 55MHz is a higher frequency, the transmission out the dark port is larger and is a closer match to the PRM transmission, so the finesse of the PRC for 55MHz is higher. 

Since the magnitudes of the fields at the reflection port are not changing significantly, our PDH signals are being created by the relative phase of something which is anti-resonant (ex. carrier or 22MHz for sideband lock) vs. something which is resonant (ex. 11MHz or 55MHz).  Since the finesse of the 55MHz signal is larger, the accumulated phase change is greater, so we expect a larger slope to our PDH signals that involve 55MHz as compared to those that use 11MHz.

If we are comparing the contributions between -11*22 and 55*22, they both include the anti-resonant 22MHz. So, the difference in the signal strengths comes directly from the difference in phase accumulation due to the variation in the dark port transmission.

EricQ had a thought, and while I have enough awake brain cells to report the thought, we're going to have to revisit it when more of our brains are awake.  In either case, the transmission through the dark port is small compared to the transmission of the ITMs, so why don't the ITMs dominate the finesse calculation, and thus are the 11MHz and 55MHz really getting that much of a difference in finesse?  To be checked out.

Attachment 1: SensMat_PRMI_24Feb2015.pdf
SensMat_PRMI_24Feb2015.pdf
  11063   Wed Feb 25 04:21:58 2015 ericqUpdateCDSSome model updates

I changed the suspension library block to acquire the SUS_[optic]_LSC_OUT channels at 16k for sensing matrix investigations. We could save the FB some load by disabling these and oplev channels in the mode cleaner optic suspensions. 

I removed nonexistant PDs from c1cal, to try and speed it up from its constantly overflowing state. It's still pretty bad, but under 60us most of the time. 

I also cleaned out the unused IPCs for simulated plant stuff from c1scx and c1sus, to get rid of red blinkeys. 

  11062   Tue Feb 24 23:39:16 2015 JenneUpdateLSCDelay line re-installed again

About 5-10 minutes ago I just put in the modulation cancellation setup according to the recipe in http://nodus.ligo.caltech.edu:8080/40m/11032

  11061   Tue Feb 24 18:54:26 2015 ericqUpdateASCSingle arm QPD ASC stability

I've lowered the UGFs for the transmission QPD servos to ~1-2Hz, and made it just an integrator. I left the arms locked with the QPD servos on for a few hours during the daytime today, and they succesfully prevented the Y arm from losing power from alignment drift for ~4 hours. Turning the servo off caused TRY to drop to ~0.6 or so. 

The X arm was only held for 2 hours or so, because after some unlock/drift event the power was below the servo trigger threshold. However, after gently nudging ETMX to get the transmission above the threshold, the servo kicked in, and brought it right back to TRX=1.0

Unfortunately, daqd was dead for much of the day, so I don't have much data to show; the trend was inferred from the wall striptool. 

It is not proven that there aren't further issues that prevent this from working with higher / more dynamic arm powers, but this is at least a point in favor of it working. 

EDIT: Here's a screenshot of the wall StripTool. Brown is TRY, blue is TRX. The downturn at the very end is me deactivating the servos. 

There is no scientific justifcation for the 0.9 threshold. Really, I should look at the noise/SNR again, now that there is some ND filtering on the QPDs. 

Attachment 1: trend.png
trend.png
  11060   Tue Feb 24 01:57:01 2015 JenneUpdateLSCModelled effect of relative modulation phase

I have used Optickle to model the effect of changing the phase between the 11MHz and 55MHz EOMs.  Also, I have looked at what modulation order is most significant (we hope it's -22*11 and -11*22).


First, so that we can compare these numbers more directly to measurements, I have included in the model the fraction of light that gets to each PD.  I'm assuming that the Faraday is about 80% transmissive, but I don't know what the true number is. Here's a cartoon showing the attenuation factors.

EDIT, 26March2015, JCD:  REFL path updated.  See elog 11172.


To model the change in relative phase between the 11MHz and 55MHz modultions, I have held the 11MHz EOM stationary, and moved the 55MHz EOM.  Since I needed an actual distance, I used a conversion factor,

\frac{\rm m}{\rm deg} = \frac{360 \ \rm deg}{55 \ \rm MHz} (\rm speed \ of \ light \ [m/s])

The sensing matrix was measured at 143Hz.  It has been corrected from mevans-meters to Newtons as the denominator.

The big thing to notice here is that the PRCL magnitude is not changing by a factor of 20.  More like a factor of 2.  BUT, I have not yet included the fact that Koji also reduced the 55MHz modulation amplitude by a factor of 3. 

As for the mysterious degeneracy of all the REFL PDs, I think we need to take a more careful measurement.  It's possible that we were seeing the real thing for REFL33, but the others don't seem to change in degeneracy with relative modulation phase. 


Why does it even matter for the 3*f1 signal what is going on with the f2 modulation?  Well, it appears that we are definitely being dominated by the 44*11 and 55*22 components.

To check this, I restricted Optickle to various orders of modulation (ex. up to second order only includes the [-22, -11, 0, 11, 22] MHz components), and plotted them all.  The change in the signal between one trace and another is the contribution from that extra modulation order.  The traces are only minutely different between orders, after the 5th order.  So, since they're all overlapping with the 5th order trace anyway, I don't plot them.

EDIT: to clarify, when I said "up to X order", I meant up to that order in 11MHz sidebands.  Optickle is applying the 11MHz and 55MHz modulations in the same way every time, but then I specify up to what order to include in the summation of different contributions to the field at a given port.  So, for the "up to 2nd order", I only include cross terms that come from [-22, -11, 0, 11, 22] MHz.  For the next order, I only include terms that come from [-33, -22, -11, 0, 11, 22, 33] MHz, etc.  So, there are no 55MHz effects when I'm only including contributions up to 2nd order (since there is a maximum cross beatnote of 44MHz), but starting with 3rd order, I do start to see signals in, for example, REFL55 and AS55, since I get terms from -22*33 and -33*22.  The first order in 55MHz (i.e. 55MHz*Carrier) only starts to show up when I calculate "up to 5th order" and above, since that includes [-55, -44, -33, -22, -11, 0, 11, 22, 33, 44, 55] MHz.


What happens if I reduce the 55MHz modulation depth by 10dB?  Since we are dominated by 55MHz-related signals, the signal at REFL 33 goes down. The maximum change we could have seen for the REFL33 PRCL signal (difference between max of blue trace and min of orange trace) is a factor of 27.


Where are we on the x-axis of these plots, and where was the maximal cancellation place that Koji found?  I need to re-check that part of the code tomorrow, to make sure that I've included all of my contributions from different components of the (field* field) matrix.

But, the moral of the story for tonight is that at least for the REFL33 signal, it's actually plausible that the optical gain went down by a factor of 20, and that the MICH and PRCL signals were degenerate.  I suspect that the total cancellation place that Koji found was somewhere around 175 degrees on the x-axis of these plots and that our nominal place is around 0 deg - around there, both the magnitude and the phase situations are possible simultaneously.

Attachment 1: ChangeInAngle_PRCLvsMICH.png
ChangeInAngle_PRCLvsMICH.png
Attachment 2: REFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.png
REFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.png
Attachment 3: REFL33_SensMat_vsModulationPhase.png
REFL33_SensMat_vsModulationPhase.png
Attachment 4: REFL33_PRCL_nominalVsLoweredGamma.png
REFL33_PRCL_nominalVsLoweredGamma.png
Attachment 5: REFL33_MICH_nominalVsLoweredGamma.png
REFL33_MICH_nominalVsLoweredGamma.png
  11059   Mon Feb 23 21:57:13 2015 KojiUpdateLSCDelay line installed again (experiment, round 1)

Last Wednesday we tried PRMI 3f modulation cancellation. Under the 3f modulation cancellation, the PRMI could not be locked
with REFL signals, and the PRCL signal was just barely sufficient to lock PRCL with help of AS55Q MICH.

- The PRCL signal level in REFL33 was reduced by factor of 20 compared with the conventional modulation setting.
=> The 3f modulation cancellation does not chage the level of 11/22MHz sidebands, it is expected that REFL33 signal
has no significant change of the signal level. But it does.  If we change the relative phase between the modulations
at 11 and 55MHz, the signal level is recovered by factor of 5. Therefore something related to 55MHz modulation
(55MHz x 22MHz, or 44MHz x 11MHz) was contributing more than -11MHzx22MHz.

- Under the 3f demodulation cancellation, MICH signal in the REFL ports were extremely weak and there was
no hope to use it for any feedback control.

- WIth the PRMI locking by REFL33->PRCL and AS55Q->MICH, the sensing matrix was measured. All of the REFL
ports however, showed extremely degenerate sensing matrix between MICH and PRCL.

This was enough confusing to us, and we didn't draw any useful information from these. Here are some ideas to
investigate what is happening in out optical and electrical system.

- One approach is to use as simple optical setup as possible to inspect our sensing systems. For example,
we may want to try PRX/PRY/XARM/YARM cavities to check the functions of the REFL diodes and qualitatively characterize
the sensing chain (Optical gain [W/m], noise level, demodulation phase) so that we can compare these with
an interferometer seinsing model.

- Another approach is to change the mdulation setting more freely and empirically try to find the characteristic
of our optical/electrical systems. e.g. change the relative modulation phase and/or 55MHz attenuation, and try to understand
how 11-22, 11-44, 22-55, 0-33 pairs are contributing the signal.

  11058   Mon Feb 23 15:04:10 2015 ZachUpdateGeneralElog restarted

The elog crashed while I was creating an entry to the Cryo log. I restarted it with the start-elog.csh script.

  11057   Mon Feb 23 08:06:59 2015 ranaUpdateASCQPD frontend code unified

Might have to also get the OL screens that go with this new code to see, but the calibrations don't go into the matrix, but rather into the OLPIT/OLYAW filter banks which follow the division, but before the servo filter banks.

  11056   Fri Feb 20 19:09:48 2015 ericqUpdateASCQPD frontend code unified

I have changed all of the oplevs and transmon QPDs to use the common ISC QPD library block, which differs mainly in its divide by zero protection. 

c1scx.mdl and c1scy.mdl were directly changed for the transmon QPDs. The oplevs were done by changing the sus_single_control.mdl library part, which is used for all of the SOSs. 

Then, because of the underscore introduced (i.e. OLPIT becomes OL_PIT because there is an OL block), I went on a sed safari to find and replace the new channel names into:

  • The filter ini files
  • various MEDM screens
  • The optic misaligning scripts (which currently live in medm/MISC/ifoalign, and need to get moved to scripts/)
  • A recent BURT snapshot, to restore all of the switches and settings easily. 
  • scripts/activateDQ.py, which is responsible for renaming OL_PIT_IN1_DQ to OPLEV_PERROR, etc.

I've fixed everything that occured to me, and the usual ways I'm used to interacting with the oplevs all seem to work at this time, but it's entirely possible I've overlooked something.

One important note is: because we are now using an effectively immutable QPD library block, the oplev urad conversion has to take place in the DoF matrix. The EPICS records C1:SUS-[OPTIC]_OL_[DOF]_CALIB still exist, but do not multiply the fast signals. Rather, the OL_MTRX elements are multiples of the CALIB value. I thought about making a new QPD_CALIBRATED part or something, but then we're right back to using custom code, which is what we're trying to avoid. 

All of the oplev DoFs are stable, I checked a few loop TFs like ETMY pitch and PRM yaw, and they looked normal. 

  11055   Fri Feb 20 14:44:47 2015 ericqUpdateCDSBad CDS behavior

In an effort to ease the IO load on the framebuilder, I've cleaned up the DQ channels being written to disk. The biggest impact was seven 64kHz channels being written to disk, on ADC channels corresponding to microphones. 

The frame files have gone from 75MB to 57MB. 

  11054   Fri Feb 20 12:29:01 2015 ericqUpdateCDSAll optics damped

I noticed some diagnostic bits in the c1sus IOP complaining about user application timing and FIFO under/overflow (The second and fourth squares next to the DACs on the GDS screen.) Over in crackle-cymac land, I've seen this correlate with large excess DAC noise. After restarting all models, all but one of these is green again, and the optics are now all damped. 

It seems there were some fishy BURT restores, as I found the TT control filters had their inputs and outputs switched off. Some ASS filters were found this way too. More undesired settings may still lurk in the mists...

The interferometer is now realigned, arms locked. 

  11053   Fri Feb 20 12:08:10 2015 ericqUpdateCDSBad CDS behavior

I've been able to get all models running. Most optics are damped, but I'm having trouble with the ITMs, BS and PRM. 

  11052   Thu Feb 19 23:23:52 2015 ChrisUpdateCDSBad CDS behavior

The frontends have some paths NFS-mounted from fb. fb is on the ragged edge of being I/O bound. I'd suggest moving those mounts to chiara. I tried increasing the number of NFS threads on fb (undoing the configuration change I'd previously made here) and it seems to help with EPICS smoothness -- although there are still occasional temporal anomalies in the time channels. The daqd flakiness (which was what led me to throttle NFS on fb in the first place) may now recur as well.

 

Quote:

At about 10AM, the C1LSC frontend stopped reporting any EPICS information. The arms were locked at the time, and remained so for some hours, until I noticed the totally whited-out MEDM screens. The machine would respond to pings, but did not respond to ssh, so we had to manually reboot.

Soon thereafter, we had a global 15min EPICS freeze, and have been in a weird state ever since. Epics has come back (and frozen again), but the fast frontends are still wonky, even when EPICS is not frozen. Intermittantly, the status blinkers and GPS time EPICS values will freeze for multiple seconds at a time, sporadically updating. Looking at a StripTool trace of an IOPs GPS time value shows a line with smooth portions for about 30 seconds, about 2 minutes apart. Between this is totally jagged step function behavior. C1LSC needed to be power cycled again; trying to restart the models is tough, because the EPICS slowdown makes it hard to hit the BURT button, as is needed for the model to start without crashing.

The DAQ network switch, and martian switch inside were power cycled, to little effect. I'm not sure how to diagnose network issues with the frontends. Using iperf, I am able to show hundreds of Mbit/s bandwidth betweem the control room machines and the frontends, but their EPICS is still totally wonky. 

What can we do??? indecision

 

  11051   Thu Feb 19 15:45:43 2015 ericqUpdateCDSBad CDS behavior

At about 10AM, the C1LSC frontend stopped reporting any EPICS information. The arms were locked at the time, and remained so for some hours, until I noticed the totally whited-out MEDM screens. The machine would respond to pings, but did not respond to ssh, so we had to manually reboot.

Soon thereafter, we had a global 15min EPICS freeze, and have been in a weird state ever since. Epics has come back (and frozen again), but the fast frontends are still wonky, even when EPICS is not frozen. Intermittantly, the status blinkers and GPS time EPICS values will freeze for multiple seconds at a time, sporadically updating. Looking at a StripTool trace of an IOPs GPS time value shows a line with smooth portions for about 30 seconds, about 2 minutes apart. Between this is totally jagged step function behavior. C1LSC needed to be power cycled again; trying to restart the models is tough, because the EPICS slowdown makes it hard to hit the BURT button, as is needed for the model to start without crashing.

The DAQ network switch, and martian switch inside were power cycled, to little effect. I'm not sure how to diagnose network issues with the frontends. Using iperf, I am able to show hundreds of Mbit/s bandwidth betweem the control room machines and the frontends, but their EPICS is still totally wonky. 

What can we do??? indecision

  11050   Thu Feb 19 04:16:45 2015 JenneUpdateLSCALS fool attempt with PRFPMI

[Jenne, EricQ]

We tried several times tonight to engage the Fool path with the PRFPMI.  No success. 

First, we locked the arms on ALS, in CARM/DARM mode, and measured the cancellation ability, to make sure that the filter shape and gain was set correctly.  For the PRFPMI, it was okay using the same shape as the single arm case, but the gain was +20.0.  There might be a bit more cancellation to be gained if we adjust the shape at the ~1dB level, but we're already able to get 20dB of cancellation, so we decided that would be enough to give things a try.  To get this much cancellation, we set the phase tracker loops to both have 2kHz UGFs, almost exactly.  We should implement a UGF servo, or the amplitude method version of that as Koji suggested ages ago, so that the phase tracker is always at the same place.

I don't think that REFL 11 is seeing as much CARM as I expect.  We ended up switching over to linearized REFL55 for our attempts.  When we're close to zero CARM offset, the arms are constantly flashing through resonance, and we get the loud buzzing.  REFL11 doesn't seem to see any of this, even though we should be close enough to see some PDH action.  REFL55 does change as we get closer to resonance, so I think it's seeing some real CARM stuff.

We tried engaging the Fool, but I don't think it did anything too useful. We need to make an estimate of what we expect our gain of the REFL loop to be - or at least the sign.

The PRMI is still not stable enough.  It keeps falling out of lock when we get to high-ish arm powers.  Not good.  More brain power tomorrow on the modulation cancellation issue. 

Perhaps if things are stable at moderate arm powers, we can use an excitation to line up the ALS vs. REFL error signals, and then watch the noises of them change as we move around in CARM offset.  This should tell us when the linearized REFL signal is quiet enough that it's worth triggering and trying to transfer over.

 

The last lockloss tonight, there was something funny going on, that we can't explain.  Even though both arms were locked on the CARM/DARM combined ALS signals, beatx doesn't see the giant oscillation that causes carm to lose lock until much later.  Fool was trying to do something, but that should affect both als individal signals in the same way.  Mystery.

Attachment 1: BEATYcrazy_duringCARMdarmLock_18Feb2015.png
BEATYcrazy_duringCARMdarmLock_18Feb2015.png
Attachment 2: 19Feb.png
19Feb.png
  11049   Thu Feb 19 04:09:21 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

X end QPD has recieved 0.2+0.4 absorptive ND filters. Y end QPD got one at 0.6. This appears to have mitigated the saturations for now; the unwhitened signals no longer go negative. The digital gains have been reset. 

 

  11048   Wed Feb 18 19:06:40 2015 KojiUpdateLSCALS Fool impulse response

ELOG V3.1.3-